OpenStudy (anonymous):

how much time has elapsed? 11:15 am 12:50pm (please show how you got that answer)it's usually at the tip of my tongue but i'm bad under preasure

OpenStudy (studygurl14):

11:15 am 12:50 pm --> 24:50 25 - 11 = 14 50 - 15 = 35 So...14 hours and 35 mins

OpenStudy (anonymous):

that doesn't make sense study girl my teacher says you have to minus the 2 not make it complicated for me! (sorry about this i'm in essential math and i'm feeling stressed because this is a portfolio and it's a big part of my grade)

OpenStudy (studygurl14):

what do you mean, minus the 2?

OpenStudy (studygurl14):

wiat, sorry i messed up.

OpenStudy (studygurl14):

do this: 11:15 am 12:50 pm 12 - 11 = 1 50 - 15 = 35 So...1 hour and 35 mins

OpenStudy (studygurl14):

Sorry, I had done 11:15 am to 12:50 am.

OpenStudy (studygurl14):

My apologies